OpenStudy (anonymous):

Ok so I understand general and slope intercept form, but it is problems like this one that get me stumped. The slope of a line is -1/5, and the y-intercept is 5. What is the equation of the line written in general form? x - 5y - 25 = 0 x - 5y - 5 = 0 x + 5y - 25 = 0 @Abhisar

OpenStudy (texaschic101):

in y = mx + b form, the slope is in the m position and the y intercept is in the b position. So if slope is -1/5 and y intercept is 5, then you have y = -1/5x + 5...but this is in slope intercept form and now we need to turn it into general form of Ax + By + C = 0. General form does not allow for fractions. y = -1/5x + 5 --- when you have fractions, you can multiply everything by the common denominator to get rid of them. 5(y) = 5(-1/5)x + 5(5) 5y = -x + 25 --- add x to both sides x + 5y = 25 --- subtract 25 from both sides x + 5y - 25 = 0 <=== General form

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

Ohh, I see. That makes sense, @texaschic101 C: Thank you!

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

Fan too? :)

OpenStudy (anonymous):

As soon as it lets me CX

OpenStudy (texaschic101):

sure...if you want to. I don't mind :) friends are good to have

OpenStudy (anonymous):

Indeed they are. Thanks again :)